OpenStudy (bethgrace96):

MEDALLLLL! solve: sqrt(x-4) +4=x

OpenStudy (anonymous):

sqrt(x-4)+4=x sqrt(x-4) =x-4 squaring bouth sides: x-4=(x-4)^2 now think. Which number is equal to itself squared?

OpenStudy (mathstudent55):

What do you get when you square a square root? For example, what is \((\sqrt{x})^2\) = ?

OpenStudy (bethgrace96):

is the answer 3,5?

OpenStudy (mathstudent55):

No.

OpenStudy (bethgrace96):

answer choices are: a)-3,4 B)3,5 C)4,5 D)4,6

OpenStudy (bethgrace96):

you make no sense dude.

OpenStudy (mathstudent55):

You can either solve the equation and find the answers, or you can substitute each pair of numbers from each choice and see for which choice both numbers work.

OpenStudy (bethgrace96):

so whats the answer

OpenStudy (mathstudent55):

Solve the equation or try the choices.

OpenStudy (anonymous):

x-4=(x-4)^2 You have a quadratic equation above from @myko's work. Solve the following for x:\[x^2-9 x+20=0 \]

OpenStudy (bethgrace96):

Thank you for making sense.

OpenStudy (bethgrace96):

so its 4,5?

OpenStudy (anonymous):

Yes. Congratulations!